OpenStudy (anonymous):

Use an equation to solve: 20% of what number is 27?

OpenStudy (anonymous):

20% of 135 is 27

OpenStudy (anonymous):

What does "of" represent? It generally means to multiply. Using that you can solve a simple equation.

OpenStudy (anonymous):

Let X=number 20% = 20/100 = 0.2 0.2X = 27 X = 27/0.2 X = 135